# Partition Functions of Determinantal Point Processes on Polarized Kähler Manifolds

### v1

@@ 1. Introduction
In this paper, we study the asymptotic expansion of the partition functions of determinantal point processes defined on a polarized Kähler manifold. These processes were introduced by Berman in a series of papers [Ber14, Ber13, Ber17] and have been shown to satisfy a large deviation principle.
We derive a full asymptotic expansion of the partition functions in two ways: one using Bergman kernel asymptotics and the other using the Quillen anomaly formula along with the asymptotic expansion of the Ray-Singer analytic torsion. By combining these two expressions, we analyze the properties of each coefficient in the asymptotic expansion. Each coefficient is given by geometric functionals on Kähler metrics that satisfy the cocycle identity, and its first variation can be expressed in terms of the coefficients of the Bergman kernel asymptotics. In particular, the second coefficient corresponds to the Mabuchi functional in Kähler geometry, while the third coefficient naturally generalizes the Liouville functional on Riemann surfaces to higher dimensions.
@@ 1.1. Mathematical background
where $\Psi^{k}$ is a holomorphic section of $(L^{k})^{\boxtimes d_{k}}$ over $M^{d_{k}}$ given by a Slater determinant. A more detailed definition will be given in Section 3. It was first introduced by Berman and subsequently studied by him, particularly regarding its relation to canonical Kähler metrics [Ber18, Ber22, Ber24]. See also [Fuj16, Aoi24] for related results. The first term in the large $k$ asymptotics of the normalization factor $Z_{k}[h]$, also known as the partition function, was obtained and proved to be equal to the Aubin-Yau functional in [BB10]. In this paper, we derive a full asymptotic expansion of $Z_{k}[h]$ as $k\rightarrow\infty$ and identify its coefficients with geometric functionals on Kähler metrics.
@@
From the Bergman kernel asymptotics approach, the first variation of the geometric functionals is expressed in terms of integrals of scalar quantities involving the curvature of the Kähler metrics, arising from the Bergman kernel asymptotic expansion. In particular, the critical point equation is of the form $a_{j}-\Delta a_{j-1}\equiv\text{constant}$, where $a_{j}$ is the $j^{th}$ coefficient in the Bergman kernel asymptotics (see 2.2). Since the works of Catlin, Ruan, Tian, and Zelditch [Cat99, Rua98, Tia90, Zel98], Bergman kernel asymptotics has become a well-studied subject with various approaches; see, for example, [DLM06, BBS08, DK10, Xu12, HKSX16]. Especially, we will use the explicit expression for $a_{j},\;j\leq 3$ that Lu computed in [Lu00]. By the result of [LT04], we know that the equation $a_{j}-\Delta a_{j-1}\equiv\text{constant}$ is an elliptic equation of order $22j+$ 2 in Kähler potential. In fact, our geometric functionals $S_{j}$ (see Theorem 3.2 for the definition) can be regarded as the functionals they sought to find in [LT04].
@@ 1.2. Relation to physics literature
Recently, higher-dimensional QHE has also been studied in cases where the even-dimensional spatial manifold has a complex structure [KN16, KN23, AKN25]. In particular, an effective action for higher-dimensional QHE was derived in terms of Chern-Simons forms. We provide an alternative derivation of this result in Section 6.
In [Kle14], based on direct computations in dimension $n=1$, Klevtsov conjectured the form of the asymptotic expansion of partition functions in all dimensions. In Section 3, we confirm his conjecture (see Remark 2). We also briefly discuss the relationship between the asymptotics of the QHE partition function and the 2D quantum gravity model in Section 5. Since this discussion is independent of the other parts of the paper, we defer it to Section 5.
@@ 1.3. Statement of the main results
Now we state our main results, starting with the asymptotic expansion of the DPP partition functions. Note that in the main text, we will use a slightly different normalization convention, which will introduce additional $2\pi$ factors. Let $\mathcal{K}_{0}$ be the space of Kähler forms in $[\omega]$ and $\mathcal{K}_{\omega}$ be the space of Kähler potentials.
@@
$S_{eff}=2i\int_{M\times[0,1]}\left[\operatorname{Td}(R_{\mathbf{T^{\prime}M}}(\bm{\omega}))CS(\operatorname{ch};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}})+CS(\operatorname{Td};\bm{\nabla}_{\mathbf{T^{\prime}M}},\bm{\nabla^{0}}_{\mathbf{T^{\prime}M}})\operatorname{ch}(R_{\mathbf{L}}(\mathbf{h_{0}}))\right]_{2n+1}+\widetilde{S}.$ (1.11)
@@
$2i\int_{M\times[0,1]}CS(\operatorname{ch}_{n+1};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}}).$ (1.12)
@@ 1.4. Structure of the paper
Acknowledgements: This research was supported by NRF Grant RS-2024-00346651.
@@ 2.1. Bergman kernel asymptotics
$\rho_{k}(\omega)=\sum_{i=1}^{d_{k}}|S_{i}^{k}|_{h^{k}}^{2}\in C^{\infty}(M,\mathbb{R}),$ (2.1)
@@ 2.2. Invariant polynomials and secondary characteristic forms
The corresponding $p$-linear functions are then defined by polarization. Note that we omit $2\pi$ factors from their standard definitions altogether so that the asymptotic Riemann-Roch-Hirzebruch formula for $L^{k}$ as $k\rightarrow\infty$ now reads (which follows from Riemann-Roch-Hirzebruch and Kodaira-Serre vanishing theorem):
@@ 2.3. Ray-Singer analytic torsion and Quillen anomaly formula
Here, the local coefficient means that it can be expressed as an integral of a density defined locally over $M$. In particular, $\beta_{j}$ is given by $-\mathcal{M}[\int_{M}\operatorname{Tr}_{s}[N{a}_{j,u}(x)]\frac{\omega^{n}}{n!}]^{\prime}(0)$ where ${a}_{j,u}$ is given by universal polynomial in $R(\omega)$ and its derivatives [DLM06, Theorem 1.2, 4.17]. Thus, if we change the Kähler form $\omega$ by the biholomorphism of $M$, $\beta_{j}$ remains the same. That is, for $f\in Aut(M)$, $\beta_{j}(\omega)=\beta_{j}(f^{*}\omega)$ where $\beta_{j}$ is understood as a formal expression. Note that from the explicit formula for $\beta_{0}$ [BV89, Theorem 8] and $\beta_{1}$ [Fin18, Theorem 1.3], we observe that $\beta_{0}$ and $\beta_{1}$ only depend on $[\omega]$. For notational convenience, set $\beta_{-1}:=0$.
3. Asymptotic expansion of the DPP Partition Functions
We start with the basic definitions regarding the determinantal point processes associated with $(L^{k},h^{k})$ for each $k\in\mathbb{N}$, with $\omega=iR(h)$ being a Kähler form. Fix a basis $(\psi_{i}^{k})_{i=1}^{d_{k}}$ of $H^{0}(M,L^{k})$. Denote by $\Psi^{k}$ the holomorphic section of $(L^{k})^{\boxtimes d_{k}}$ over $M^{d_{k}}$ defined by a Slater determinant
@@ 4. Holomorphic Invariants from Bergman Kernel Asymptotics and Critical Points
We will show that $\frac{d}{dt}F(\omega_{t},X)=0$. Let $\theta_{t}$ be a potential function of $X$ with respect to $\omega_{t}$, that is, $\iota_{X}\omega_{t}=-\bar{\partial}\theta_{t}$. Setting $\alpha_{t}:=-i\partial\dot{\varphi_{t}}$, we have $\dot{\omega_{t}}=i\partial\bar{\partial}\dot{\varphi_{t}}=\bar{\partial}\alpha_{t}$ and
@@
where $\Gamma_{t}$ denotes the Levi–Civita connection $1$-form for $\omega_{t}$. Note that $\dot{\Gamma_{t}}$ is globally well defined and $\dot{R_{t}}=\bar{\partial}\dot{\Gamma_{t}}$. Using these, we compute $\frac{d}{dt}F(\omega_{t},X)$ (we suppress the subscript $t$):
@@ 6. Derivation of the (2n+1)D Chern-Simons action
$Z[A_{\mu}]=e^{-S_{eff}[A_{\mu}]},$ (6.1)
in Euclidean signature. The functional derivative of the effective action with respect to the time component of $A_{\mu}$ is given by
@@
Let $\varphi_{1}\in\mathcal{K}_{\omega}$ and choose a smooth path $\varphi_{t}$ in $\mathcal{K}_{\omega}$ joining 00 and $\varphi_{1}$. Let $\varphi:\mathbb{C}\rightarrow\mathcal{K}_{\omega}$ be a smooth map defined by $\varphi(z)=\varphi_{t}$ where $z=t+is\in\mathbb{C}$ and trivially extended over $t\notin[0,1]$. Define Hermitian metrics $\mathbf{h}$ and $\bm{\omega}$ on the pull-back bundles $\mathbf{L}:=\pi_{1}^{*}L$ and $\mathbf{T^{\prime}M}:=\pi_{1}^{*}T^{1,0}M$, where $\pi_{1}:M\times\mathbb{C}\rightarrow M$ is the projection, by
@@
$S_{eff}[A_{\mu}]=-\log Z_{1}[\varphi_{1}]$ (6.13)
$=i\int_{M}BC(\operatorname{Td};\omega_{\varphi_{1}},\omega)\operatorname{ch}(R_{L}(h))+\operatorname{Td}(R_{T^{1,0}M}(\omega_{\varphi_{1}}))BC(\operatorname{ch};h_{\varphi_{1}},h)-2\log\frac{T(\omega_{\varphi_{1}},h_{\varphi_{1}})}{T(\omega,h)}.$ (6.14)
Denote $\widetilde{S}:=-2\log\frac{T(\omega_{\varphi_{1}},h_{\varphi_{1}})}{T(\omega,h)}$. Now we derive [KN16, (26), (41)].
@@
$S_{eff}=2i\int_{M\times[0,1]}\left[\operatorname{Td}(R_{\mathbf{T^{\prime}M}}(\bm{\omega}))CS(\operatorname{ch};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}})+CS(\operatorname{Td};\bm{\nabla}_{\mathbf{T^{\prime}M}},\bm{\nabla^{0}}_{\mathbf{T^{\prime}M}})\operatorname{ch}(R_{\mathbf{L}}(\mathbf{h_{0}}))\right]_{2n+1}+\widetilde{S}.$ (6.15)
@@
$2i\int_{M\times[0,1]}CS(\operatorname{ch}_{n+1};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}}).$ (6.16)
@@
$\int_{M}BC(\operatorname{Td};\omega_{\varphi_{1}},\omega)\operatorname{ch}(R_{L}(h))+\operatorname{Td}(R_{T^{1,0}M}(\omega_{\varphi_{1}}))BC(\operatorname{ch};h_{\varphi_{1}},h)$ (6.17)
$=\int_{0}^{1}\frac{\partial}{\partial t}\left(\int_{M}BC(\operatorname{Td};\omega_{\varphi_{t}},\omega)\operatorname{ch}(R_{L}(h))+\operatorname{Td}(R_{T^{1,0}M}(\omega_{\varphi_{t}}))BC(\operatorname{ch};h_{\varphi_{t}},h)\right)\wedge dt$ (6.18)
@@ Appendix A Explicit Computations on the Generalized Liouville Action
$\psi_{1}\cdot\gamma^{2}(\psi_{2})=\left(\text{terms symmetric in $\psi_{1}$ and $\psi_{2}$}\right)$ (A.39)

### v2

@@ 1. Introduction
In this paper, we study the asymptotic expansion of the partition functions of determinantal point processes defined on a polarized Kähler manifold. These processes were introduced by Berman in a series of papers [Ber13, Ber14, Ber17] and have been shown to satisfy a large deviation principle.
We derive a full asymptotic expansion of the partition function in two ways: one using Bergman kernel asymptotics and the other using the Quillen anomaly formula along with the asymptotic expansion of the Ray-Singer analytic torsion. By combining these two expressions, we analyze the properties of each coefficient in the asymptotic expansion. Each coefficient is given by geometric functionals on Kähler metrics that satisfy the cocycle identity, and its first variation can be expressed in terms of the coefficients of the Bergman kernel asymptotics. In particular, the second coefficient corresponds to the Mabuchi functional in Kähler geometry, while the third coefficient naturally generalizes the Liouville functional on Riemann surfaces to higher dimensions.
@@ 1.1. Mathematical background
where $\Psi^{k}$ is a holomorphic section of $(L^{k})^{\boxtimes d_{k}}$ over $M^{d_{k}}$ given by the Slater determinant. A more detailed definition will be given in Section 3. It was first introduced by Berman and subsequently studied by him, particularly regarding its relation to canonical Kähler metrics [Ber18, Ber22, Ber24]. See also [Fuj16, FO18, Aoi24] for related results.
The first term in the large $k$ asymptotics of the normalization factor $Z_{k}[h]$, also known as the partition function, was obtained and proved to be equal to the Aubin-Yau functional by Berman-Boucksom [BB10]. In fact, they considered a more general setting, and their results have been recently extended in [Fin25]. In this paper, we focus on the (smooth) polarized case and derive the full asymptotic expansion of $Z_{k}[h]$ as $k\rightarrow\infty$, identifying its coefficients with geometric functionals on Kähler metrics.
@@
From the Bergman kernel asymptotics approach, the first variation of the geometric functionals is expressed in terms of integrals of scalar quantities involving the curvature of the Kähler metrics, arising from the Bergman kernel asymptotic expansion. In particular, the critical point equation is of the form $a_{j}-\Delta a_{j-1}\equiv\text{constant}$, where $a_{j}$ is the $j^{th}$ coefficient in the Bergman kernel asymptotics; see (2.2). Since the works of Catlin, Ruan, Tian, and Zelditch [Cat99, Rua98, Tia90, Zel98], Bergman kernel asymptotics has become a well-studied subject with various approaches; see, for example, [DLM06, BBS08, DK10, Xu12, HKSX16]. Especially, we will use the explicit expression for $a_{j},\;j\leq 3$ that Lu computed in [Lu00]. By the result of [LT04], we know that the equation $a_{j}-\Delta a_{j-1}\equiv\text{constant}$ is an elliptic equation of order $22j+$ 2 in Kähler potential. In fact, our geometric functionals $S_{j}$ (see Theorem 3.2 for the definition) can be regarded as the functionals they sought to find in [LT04].
@@ 1.2. Relation to physics literature
Recently, higher-dimensional QHE has also been studied in cases where the even-dimensional spatial manifold has a complex structure [KN16, KN23, AKN25]. In particular, an effective action for higher-dimensional QHE was derived in terms of Chern-Simons forms. We provide an alternative derivation of this result in Section 6. Note that in dimension $n=1$, [KMMW17] derived the effective action in terms of the Chern-Simons functional considering a more general moduli space.
In [Kle14], based on direct computations in dimension $n=1$, Klevtsov conjectured the form of the asymptotic expansion of partition function in all dimensions. We confirm his conjecture in Section 3 (see Remark 3). We also briefly discuss the relationship between the asymptotics of the QHE partition function and the 2D quantum gravity model in Section 5. Since this discussion is independent of the other parts of the paper, we defer it to Section 5.
@@ 1.3. Statement of the main results
Now we state our main results, starting with the asymptotic expansion of the DPP partition function. Note that in the main text, we will use a slightly different normalization convention, which will introduce additional $2\pi$ factors. Let $\mathcal{K}_{0}$ be the space of Kähler forms in $[\omega]$ and $\mathcal{K}_{\omega}$ be the space of Kähler potentials.
@@
$S_{eff}=2\int_{M\times[0,1]}\left[\operatorname{Td}(R_{\mathbf{T^{\prime}M}}(\bm{\omega}))CS(\operatorname{ch};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}})+CS(\operatorname{Td};\bm{\nabla}_{\mathbf{T^{\prime}M}},\bm{\nabla^{0}}_{\mathbf{T^{\prime}M}})\operatorname{ch}(R_{\mathbf{L}}(\mathbf{h_{0}}))\right]_{2n+1}+\widetilde{S}.$ (1.11)
@@
$2\int_{M\times[0,1]}CS(\operatorname{ch}_{n+1};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}}).$ (1.12)
@@ 1.4. Structure of the paper
Acknowledgements: The author would like to thank Siarhei Finski for clarifications regarding his Theorem and Remark 1. This research was supported by NRF Grant RS-2024-00346651.
@@ 2.1. Bergman kernel asymptotics
$\rho_{k}(\omega)=\sum_{i=1}^{d_{k}}||S_{i}^{k}||_{h^{k}}^{2}\in C^{\infty}(M,\mathbb{R}),$ (2.1)
@@ 2.2. Invariant polynomials and secondary characteristic forms
The corresponding $p$-linear functions are then defined by polarization. Note that we omit $2\pi$ factors from their standard definitions altogether so that the asymptotic Riemann-Roch-Hirzebruch formula for $L^{k}$ as $k\rightarrow\infty$ now reads (which follows from Riemann-Roch-Hirzebruch theorem and Kodaira-Serre vanishing theorem):
@@ 2.3. Ray-Singer analytic torsion and Quillen anomaly formula
Here, the local coefficient means that it can be expressed as an integral of a density defined locally over $M$. More precisely, $\beta_{j}$ is given by $-\mathcal{M}[\int_{M}\operatorname{Tr}_{s}[N{a}_{j,u}(x)]\frac{\omega^{n}}{n!}]^{\prime}(0)$ where ${a}_{j,u}$ is given by universal polynomial in $R(\omega)$ and its derivatives [DLM06, Theorem 1.2, 4.17]. Thus, if we change the Kähler form $\omega$ by the biholomorphism of $M$, $\beta_{j}$ remains the same. That is, for $f\in Aut(M)$, $\beta_{j}(\omega)=\beta_{j}(f^{*}\omega)$ where $\beta_{j}$ is understood as a formal expression. Note that from the explicit formula for $\beta_{0}$ [BV89, Theorem 8] and $\beta_{1}$ [Fin18, Theorem 1.3], we observe that $\beta_{0}$ and $\beta_{1}$ only depend on $[\omega]$. For notational convenience, set $\beta_{-1}:=0$.
Remark 1.
Indeed for $f\in Aut(M)$, one can verify that $\square_{L^{k}}=(f^{-1})^{*}\circ\square_{f^{*}L^{k}}\circ f^{*}$, where $\square_{f^{*}L^{k}}$ is defined with respect to $f^{*}\omega$ and $f^{*}h^{k}$. Thus, the spectra of the two Laplacians, and consequently the analytic torsions corresponding to $\omega$ and $f^{*}\omega$, coincide. Since both analytic torsions admit asymptotic expansions by Theorem 2.2, the coefficients of these expansions, $\beta_{j}(\omega)$ and $\beta_{j}(f^{*}\omega)$, must be identical. The point of the above argument using the local nature of $\beta_{j}$ is that it does not involve the polarizing line bundle $L$.
3. Asymptotic expansion of the DPP partition function
We start with the basic definitions regarding the determinantal point processes associated with $(L^{k},h^{k})$ for each $k\in\mathbb{N}$, with $\omega=iR(h)$ being a Kähler form. Fix a basis $(\psi_{i}^{k})_{i=1}^{d_{k}}$ of $H^{0}(M,L^{k})$. Denote by $\Psi^{k}$ the holomorphic section of $(L^{k})^{\boxtimes d_{k}}$ over $M^{d_{k}}$ defined by the Slater determinant
@@ 4. Holomorphic invariants from Bergman kernel asymptotics and critical points
We will show that $\frac{d}{dt}\widetilde{F_{j}}(\omega_{t},X)=0$. Let $\theta_{t}$ be a potential function of $X$ with respect to $\omega_{t}$, that is, $\iota_{X}\omega_{t}=-\bar{\partial}\theta_{t}$. Setting $\alpha_{t}:=-i\partial\dot{\varphi_{t}}$, we have $\dot{\omega_{t}}=i\partial\bar{\partial}\dot{\varphi_{t}}=\bar{\partial}\alpha_{t}$ and
@@
where $\Gamma_{t}$ denotes the Levi–Civita connection $1$-form for $\omega_{t}$. Note that $\dot{\Gamma_{t}}$ is globally well defined and $\dot{R_{t}}=\bar{\partial}\dot{\Gamma_{t}}$. Using these, we compute $\frac{d}{dt}\widetilde{F_{j}}(\omega_{t},X)$ (we suppress the subscript $t$):
@@ 6. Derivation of the (2n+1)D Chern-Simons action
$Z[A_{\mu}]=e^{iS_{eff}[A_{\mu}]}.$ (6.1)
The functional derivative of the effective action with respect to the time component of $A_{\mu}$ is given by
@@
Let $\varphi_{1}\in\mathcal{K}_{\omega}$ and choose a smooth path $\varphi_{t}$ in $\mathcal{K}_{\omega}$ joining 00 and $\varphi_{1}$. Let $\varphi:\mathbb{C}\rightarrow\mathcal{K}_{\omega}$ be a smooth map defined by $\varphi(z)=\varphi_{t}$ where $z=1-t+is\in\mathbb{C}$ and trivially extended over $t\notin[0,1]$. Define Hermitian metrics $\mathbf{h}$ and $\bm{\omega}$ on the pull-back bundles $\mathbf{L}:=\pi_{1}^{*}L$ and $\mathbf{T^{\prime}M}:=\pi_{1}^{*}T^{1,0}M$, where $\pi_{1}:M\times\mathbb{C}\rightarrow M$ is the projection, by
@@
$S_{eff}[A_{\mu}]=-i\log Z_{1}[\varphi_{1}]$ (6.13)
$=-\int_{M}BC(\operatorname{Td};\omega_{\varphi_{1}},\omega)\operatorname{ch}(R_{L}(h))+\operatorname{Td}(R_{T^{1,0}M}(\omega_{\varphi_{1}}))BC(\operatorname{ch};h_{\varphi_{1}},h)-2i\log\frac{T(\omega_{\varphi_{1}},h_{\varphi_{1}})}{T(\omega,h)}.$ (6.14)
Denote $\widetilde{S}:=-2i\log\frac{T(\omega_{\varphi_{1}},h_{\varphi_{1}})}{T(\omega,h)}$. Now we derive [KN16, (26), (41)].
@@
$S_{eff}=2\int_{M\times[0,1]}\left[\operatorname{Td}(R_{\mathbf{T^{\prime}M}}(\bm{\omega}))CS(\operatorname{ch};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}})+CS(\operatorname{Td};\bm{\nabla}_{\mathbf{T^{\prime}M}},\bm{\nabla^{0}}_{\mathbf{T^{\prime}M}})\operatorname{ch}(R_{\mathbf{L}}(\mathbf{h_{0}}))\right]_{2n+1}+\widetilde{S}.$ (6.15)
@@
$2\int_{M\times[0,1]}CS(\operatorname{ch}_{n+1};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}}).$ (6.16)
@@
$-\int_{M}BC(\operatorname{Td};\omega_{\varphi_{1}},\omega)\operatorname{ch}(R_{L}(h))+\operatorname{Td}(R_{T^{1,0}M}(\omega_{\varphi_{1}}))BC(\operatorname{ch};h_{\varphi_{1}},h)$ (6.17)
$=-\int_{0}^{1}\frac{\partial}{\partial t}\left(\int_{M}BC(\operatorname{Td};\omega_{\varphi_{t}},\omega)\operatorname{ch}(R_{L}(h))+\operatorname{Td}(R_{T^{1,0}M}(\omega_{\varphi_{t}}))BC(\operatorname{ch};h_{\varphi_{t}},h)\right)\wedge dt$ (6.18)
@@ Appendix A Explicit computations on the generalized Liouville action
$\psi_{1}\cdot\gamma^{2}(\psi_{2})=\left(\text{terms symmetric in $\psi_{1}$ and $\psi_{2}$}\right)$